OpenStudy (anonymous):

let x to the power m+y to the power n =1 if dy/dx= -y to the power 1/3 .then what is the value of m?

OpenStudy (anonymous):

@myininaya @ajprincess @lgbasallote

OpenStudy (shubhamsrg):

\[x ^{(m+y)^{n}} = 1 ? \]

OpenStudy (shubhamsrg):

is it like this ?

OpenStudy (anonymous):

i can't view what you 've posted

OpenStudy (shubhamsrg):

press "ctrl" and " + " sign together,, the view maximizes

OpenStudy (shubhamsrg):

and to bring back to normal, crtl and - together..

OpenStudy (anonymous):

yah ...i saw no its not like tht.

OpenStudy (shubhamsrg):

ohh i see..is it x^m + y^n = 1 ?

OpenStudy (anonymous):

yes..

OpenStudy (shubhamsrg):

kk.. what did you try on this ? i mean where are you stuck ?

OpenStudy (shubhamsrg):

seems like differentiating both sides directly should do it probably ?

OpenStudy (anonymous):

dy/dx=-y^{1/3} y^{-1/3}dy=-dx ∫y^{-1/3}dy=-∫dx

OpenStudy (anonymous):

3y^{1/3}=-x+c now what more to do ?

OpenStudy (shubhamsrg):

please recheck your LHS,,there;s a mistake.. also,,i dont think there's any need for integration here..why dont you just differentiate the given eqn ?

OpenStudy (anonymous):

there are options.. do i post them ?

OpenStudy (shubhamsrg):

no need,,arent you following what am trying to say ?